You have plenty of cash to invest. You are considering an investment of $125,000 in a...

  1. You have plenty of cash to invest. You are considering an investment of $125,000 in a project which is expected to earn $14425 a year for ten years. Is it an attractive investment if your minimum expected annual rate of return is 5% (compound interest)? Calculate the expected annual rate of return (R) of the project. (6 points) (NO TABLES ONLY HAND CALCULATION) As well explain which equation you have used
Answer #1

As per the information given in the question

Initial investment of the project (I) =$125000

Annual expected income (A) =14425

Effective lift of the project (N) =10 years

Minimum expected annual rate of return or MARR =5%

Annual expected rate of return (ROR)=?

To calculate ROR of the project we have followed trial and error method

Present Value factor (PVF) = (1+i)-N

Years (N)

Cash flow

PVF@2%

PV of Cashflow

PVF@3%

PV of Cashflow

0

-125000

1

-125000

1

-125000.00

1

14425

0.980392

14142.16

0.970874

14004.85

2

14425

0.961169

13864.86

0.942596

13596.95

3

14425

0.942322

13593.00

0.915142

13200.92

4

14425

0.923845

13326.47

0.888487

12816.43

5

14425

0.905731

13065.17

0.862609

12443.13

6

14425

0.887971

12808.99

0.837484

12080.71

7

14425

0.870560

12557.83

0.813092

11728.85

8

14425

0.853490

12311.60

0.789409

11387.23

9

14425

0.836755

12070.19

0.766417

11055.56

10

14425

0.820348

11833.52

0.744094

10733.55

NPV at 2%

4573.79

NPV at 3%=

-1951.82

Rate of return (ROR) is the interest rate at which the NPV=0, to estimate it we have followed trial and error method, first we have taken the PV factor = 2%, where the NPV =$4573.79, which is positive and then we have taken the PV factor = 3%, where the NPV = -$1951.82 which is negative. So the Rate of return must fall in between 2% to 3%.

Therefore

R1 = 2%                               NPV1=$4573.79

R2= 3%                 NPV2=-$1951.82

Using linear interpolation method of estimation of Rate of Return

ROR= R1 + (R2-R1)X[NPV1/(NPV1-NPV2)]

ROR = 2 + (3-2)[(4573.79)/ {4573.79-(-1951.82)}]

ROR = 2 + (1)[(4573.79)/(4573.79+1951.82)]

ROR = 2 + (1)[(4573.79)/(6525.61)]=2+0.7009 =2.7009   or Approx 2.7%

The annual rate of return is approx 2.7%

Since ROR =2.7% < MARR(5%), it is not a attractive project for investment